what is the dimensions density

Grade:12th Pass

5 Answers

TANAYRAJ SINGH CHOUHAN
11 years ago

Density = Mass/Volume

so unit of density = kg/m3 

and dimension of density = [M]1[L]3

Approve if you like the answer 

Best of luck !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

rahul thengadi
11 years ago

you are wrong ,correct ans is [M1,L-3]


Raunak Sudeep Taysheti
11 years ago

density is mass per unit volume

 

therefore, density = mass/volume

                          =  M/L^3

                          ML^-3

 

MUKESH chinna YADAV
11 years ago

GD EVNG FRND YOUR ANSWER I SOLVE BELOW

 

DENSITY:- 1. the amount per unit size
 2. the spatial property of being crowded together

density=mass/volume

=kilogram/(metre)^3

=kg/m^3

Dimensions of density:-[M][L]^-3

ALL THE BEST AND STUDY WELL

G Ashwin -
11 years ago

density is mass/volume

d=M/L^3

d=ML^-3